satellite73 (satellite73):

\[\frac{4^{\left(\frac{1}{\log_4(\frac{3}{4})}\right)}}{3^{\left(\frac{1}{\log_3(\frac{3}{4})}\right)}}\]

Parth (parthkohli):

Wasnt that log4^3/4 and log_3^3/4 asked in a question before?

OpenStudy (anonymous):

it was ambiguous before apparently it is \(\frac{1}{12}\)

OpenStudy (anonymous):

tried changing the base, and also rewriting the logs as a difference, but it is not coming snappy at all

Parth (parthkohli):

So we just have to find \(\Large {4^{1 } \over 3^{1 \over 12}}\)? I'm new with logs so mercy

OpenStudy (anonymous):

having the damndest time typesetting as well

OpenStudy (anonymous):

For my own reason: \[ \huge \frac{4^{\left(\frac{1}{\log_4(\frac{3}{4})}\right)}}{3^{\left(\frac{1}{\log_3(\frac{3}{4})}\right)}} \]

OpenStudy (anonymous):

Is this right?

OpenStudy (anonymous):

yes that is it

OpenStudy (anonymous):

@experimentX you lost me there

OpenStudy (experimentx):

i got \[ \frac 4 3 \log_4 3\] is 1/12 supposed to be answer??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1/12 is the right answer.

OpenStudy (anonymous):

\[\huge \frac{4^{\left(\frac{1}{\log_4(\frac{3}{4})}\right)}}{3^{\left(\frac{1}{\log_3(\frac{3}{4})}\right)}}\] yes it is i do not see the gimmick change base? rewrite as \[\huge \frac{4^{\left(\frac{\ln(4)}{\ln(\frac{3}{4})}\right)}}{3^{\left(\frac{\ln(3)}{\ln(\frac{3}{4})}\right)}}\]

OpenStudy (anonymous):

I am gonna M.SE it .. if you don't have any objection?

OpenStudy (anonymous):

i do not know what that means, but i have no objection this was a previous problem so go nuts

OpenStudy (anonymous):

not times, to the power of

OpenStudy (anonymous):

somehow we are supposed to arrive at \(\frac{1}{4\times 3}\)

OpenStudy (experimentx):

Oh .. my mistake

OpenStudy (anonymous):

thing maybe rewrite exponents as \(\frac{\log_4(4)}{\log_4(\frac{3}{4})}\)

OpenStudy (anonymous):

http://math.stackexchange.com/questions/161685/

OpenStudy (experimentx):

i wish i could answer ... lol

OpenStudy (anonymous):

typesetting is a mess :/

OpenStudy (anonymous):

oh maybe we can change the base to \(\frac{4}{3}\)!!

OpenStudy (anonymous):

?

OpenStudy (experimentx):

it seems that i can wrap latex with $ here too .. i thought \[ only works here

OpenStudy (anonymous):

doesn't work for me $\frac{a}{b}$

OpenStudy (anonymous):

for in line i use \(

OpenStudy (anonymous):

Work : $$ \frac{a}{b}$$

OpenStudy (anonymous):

ok how about changing the base to \(\frac{4}{3}\) maybe that would work

OpenStudy (anonymous):

$$ one ;)

OpenStudy (zarkon):

I changed it to 12 to some power...then converted the exponent to natural logs...simplified and got -1 ie \[12^{-1}\] though I'm lazy and am not going to type that much

OpenStudy (anonymous):

12 to some power?

OpenStudy (zarkon):

\[12^{\log_{12}(\cdot)}\]

OpenStudy (zarkon):

the dot it the original problem

OpenStudy (zarkon):

*is the..

OpenStudy (anonymous):

you mean you converted \(\frac{4}{3}\) ... oh ok but that gimmick works if you know the answer i guess, otherwise where would the 12 come from?

OpenStudy (zarkon):

I also did it for \[a\times b\]

OpenStudy (anonymous):

christ no wonder you are not going to write it here, i can barely write it on paper

OpenStudy (zarkon):

\[\huge \frac{a^{\left(\frac{1}{\log_a(\frac{b}{a})}\right)}}{b^{\left(\frac{1}{\log_b(\frac{b}{a})}\right)}}=\frac{1}{a\times b}\]

OpenStudy (anonymous):

and why? this of course gives it right away

OpenStudy (zarkon):

i derived it...seemed a natural thing to do...combine into one exponent

OpenStudy (foolaroundmath):

\[y = \frac{a^{\frac{1}{\log_{a}(\frac{b}{a})}}}{b^{\frac{1}{\log_{3}(\frac{b}{a})}}}\] take log on both sides with base (b/a) \[\log_{b/a}y = \frac{1}{\log_{a}(b/a)}\log_{b/a}a - \frac{1}{\log_{b}(b/a)}\log_{b/a}b \] \[\log_{b/a}y =\frac{loga.loga}{\log(b/a)\log(b/a)} - \frac{logb.logb}{\log(b/a).\log(b/a)}\] \[\log_{b/a}y = \frac{(loga-logb)(loga+logb)}{\log(b/a).\log(b/a)}\] \[\log_{b/a}y = \frac{-\log(ab)}{\log(b/a)}\] \[\log_{b/a}y = \log_{b/a}(\frac{1}{ab})\] \[=> y = \frac{1}{ab}\]

OpenStudy (experimentx):

$$ \huge a = b^{\log_ba} $$ $$ \huge \frac{\log_ba}{\log_a b - 1 } - \frac{1}{1 - \log_ba} = \frac{(\log_ba)^2}{1-\log_ba} - \frac{1}{1-\log_ba} \\ \frac{-(1 + \log_ba)(1 - \log_ba)}{(1 - \log_ba)} = -(\log_b + \log_a) = -\log_b(ab) $$

OpenStudy (experimentx):

$$ \huge b^{-\log_b(ab)} = 1/ab$$

OpenStudy (anonymous):

ok here is my effort \[\large \exp{\frac{\ln(a)}{\log_a(\frac{b}{a})}}=\exp{\frac{\ln^2(a)}{\ln(b)-\ln(a)}}\] and similarly \[\large \exp{\frac{\ln(b)}{\log_b(\frac{b}{a})}}=\exp\frac{\ln^2(b)}{\ln(b)-\ln(a)}\] subtract and get \[\frac{\ln^2(a)-\ln^2(b)}{\ln(b)-\ln(a)}=-(\ln(b)+\ln(a))=-\ln(ab)\]and finally \[\exp^{-\ln(ab)}=\frac{1}{ab}\]

OpenStudy (anonymous):

using the definition \(a^x=e^{x\ln(a)}\) i have to say other ways were snappier though

OpenStudy (maheshmeghwal9):

Hmm............................ gr8* :D